How would I be able to write an equation in slope-intercept form of the line passing through two given points?
(0,7) (1,9)
(0,3) (2,-1)
(5,2) (7,0)
Thank you!

Answer by Cromlix(4381)

Hi there,
Gradient = y2 - y1/x2 - x1
Gradient = 9 - 7/1 - 0
Gradient = 2/1
Gradient = 2
Using line equation:
y - b = m(x - a)
m = 2 (a,b) = (1,9)
y - 9 = 2(x - 1)
y - 9 = 2x - 2
y = 2x - 2 + 9
y = 2x + 7
.............
Gradient = y2 - y1/x2 - x1
Gradient = -1 -3/2 - 0
Gradient = -4/2
Gradient = -2
Using line equation:
y - b = m(x - a)
m = -2 (a,b) = (2,-1)
y - (-1) = -2(x - 2)
y + 1 = -2x + 4
y = -2x + 4 - 1
y = -2x + 3
.................
Gradient = y2 - y1/x2 - x1
Gradient = 0 - 2/7 - 5
Gradient = -2/2
Gradient = -1
Using line equation:
y - b = m(x - a)
m = -1 (a,b) = (5,2)
y - 2 = -1(x - 5)
y - 2 = -x + 5
y = -x + 5 + 2
y = -x + 7
..................
Hope this helps :-)